OpenStudy (anonymous):

If 37.4 grams of water decomposes at 297 Kelvin and 1.30 atmospheres, how many liters of oxygen gas can be produced? Show all of the work used to solve this problem. 2 H2O (l) 2 H2 (g) + O2 (g)

OpenStudy (anonymous):

Let's solve this using molar analysis. First, we need to find the number of moles in 37.4 grams of water. The molecular weight of water is 36 grams per mole. Therefore, \[n = {m \over M}\]where n is the number of moles, m is the mass in grams, and M is the molecular weight in grams/mole. Let's realize that 2 moles of water creates 1 mole of oxygen. We can find the number of moles of oxygen produced by 37.4 grams of water by the following expression: \[n_{O_2} = {\rm 1 mole ~O_2 \over 2 mole~H_2O} \cdot n_{H_2O} = {\rm 1 mole ~O_2 \over 2 mole~H_2O} \cdot \left( m_{H_2O} \over M_{H_2O} \right)\] Once we know the number of moles of oxygen that 37.4 grams of water can produce, we can use the ideal gas law to find the volume of the gas. \[PV = n RT\]Use\[R =0.08205746 \left[ \rm L \cdot atm \over K \cdot mole \right]\]

OpenStudy (anonymous):

Chemistry? This is Physics. Just letting you know you'd be more successful in the Chemistry forum...